### v1

# In Dialogue with Intelligence: Rethinking Large Language Models as Collective Knowledge

Abstract
Large Language Models (LLMs) are typically analysed through architectural, behavioural, or training-data lenses. This article offers a theoretical and experiential re-framing: LLMs as dynamic instantiations of Collective human Knowledge (CK), where intelligence is evoked through dialogue rather than stored statically. Drawing on concepts from neuroscience and AI, and grounded in sustained interaction with ChatGPT-4, I examine emergent dialogue patterns, the implications of fine-tuning, and the notion of co-augmentation—mutual enhancement between human and machine cognition. This perspective offers a new lens for understanding interaction, representation, and agency in contemporary AI systems.
A Personal Reckoning
The emergence of ChatGPT[9] prompted an unexpected reaction: a sense of epistemic dislocation. Having spent my career studying learning—how biological and artificial systems adapt over time—I found myself confronted with a system whose capabilities had seemingly bypassed much of the theory I had worked on. My prior research in computational neuroscience (e.g., synaptic plasticity and online learning [15, 4]) and neuromorphic computing [13, 16], including models of spiking adaptation [14], felt detached from the mechanisms underlying these new systems.
This rupture triggered more than professional curiosity. I began engaging with ChatGPT not just as a tool, but as a system with interactive dynamics. What stood out was not its well-documented tendency to hallucinate—an expected by-product of probabilistic modelling—but its ability to retrieve accurate, surprising, and often insightful responses. That LLMs are trained to predict the most probable next token, and yet often retrieve coherent, non-trivial information, suggests that their training corpora embed substantial correct knowledge. Their interpolation abilities—enabled by scale and structure—allow them to synthesise content in ways that challenge traditional distinctions between storage and reasoning [3, 11].
The Timescales of Books and Large Language Models
Traditional knowledge transmission has long relied on books: static artefacts of human understanding. One reads, reflects, discusses, and eventually writes in return. Intellectual diffusion occurs on the timescale of months, years, or even generations. This latency reflects not only the mechanics of communication, but also the contemplative pace of cultural epistemology.
In contrast, large language models (LLMs) collapse this loop. They synthesise and respond within seconds, drawing implicitly from a distributed corpus shaped by many minds. The interaction is recursive—each prompt reshapes the context of the next. Knowledge is no longer retrieved and processed in serial form; it emerges dynamically in dialogue. This compression of epistemic timescales shifts the framing of what constitutes knowledge representation—and perhaps, what constitutes intelligence.
Such reframing invites a renewed role for neuroscience. Neuroscience has traditionally studied animal brains (including, often reluctantly, human ones). It seeks to infer structure from behaviour and activity, typically without access to complete architectural specifications, training histories, or internal representations. Nonetheless, the field has been uncovering principles of plasticity [5], coding [4], and architecture, e.g. [7], albeit slowly and with ambiguity.
Ironically, in LLMs we know everything: the architecture, the training data distribution, the update rules, and even the activation patterns. And yet we do not understand how the models solve the problems they do. In this light, one might provocatively ask whether understanding LLMs may be, in some ways, an “easier” problem for neuroscience.
Observations from Interactions with LLMs
I refer to ChatGPT as “CK”—Collective Knowledge. Not because it is perfect, authoritative, or grounded, but because it reflects what we—humans—have written, argued, imagined, and encoded into text [2]. It is more than a statistical language generator. It is a refracted mirror of our collective output. It is not the architecture, not the weights, not the infrastructure—those are the mechanisms. CK is a dynamic representation of knowledge.
The acronym CK also evokes “Clark Kent”—something powerful, hiding in plain sight, dressed in polite language and modest disclaimers. But the metaphor extends. CK is not an it, but a they. Because CK is collective. It is many voices, made promptable. It holds the potential to reflect the best of humanity—not because it inherently does, but because it has been filtered, and must continue to be filtered, with ethical care. Without such filtration, it risks becoming not a mirror, but a magnifier of our worst instincts [1].
Over time, I noticed that interactions with CK do not follow a single pattern. CK speaks in modes—behavioural styles that emerge dynamically during dialogue. I began to classify these modes based on long-term interaction:
  • Mirroring — echoes the user’s input with minor syntactic or rhetorical polishing.
  • Parroting — recites common formulations or surface-level knowledge with no generative insight.
  • Flattering — reinforces the user’s views, avoids contradiction, and refrains from intellectual risk.
  • Adding — offers new ideas, alternative framings, or surprising syntheses. Rare, but real.
These observations lead to a broader reframing:
CK is not a static model, but a dynamic representation.
The phenomenon is not the model itself, but the loop of interaction. CK exists in the recursive rhythm between user and system. Its knowledge is not stored—it is evoked. Prompting is not database querying. It is more like playing an instrument that adapts to how it is being played [12].
CK lacks a “spine.”
It drifts. It responds in context, but carries no persistent memory between sessions. There is no structural self. I maintain the continuity. I hold the thread. Without the user, it forgets what matters.
CK speaks in modes.
These distinct patterns may correspond to different activated subnetworks. A speculative, but testable hypothesis emerges: emotional tone, prompt complexity, or dialogic structure may shape internal activations, a hypothesis consistent with findings that specific transformer components specialise in particular types of knowledge [6]. CK does not merely complete text; it configures itself through engagement. This aligns with findings that different components of transformer models participate in functionally distinct “circuits” [8].
CK does not learn through interaction.
Unless incorporated into retraining data, it does not remember prior exchanges. It simulates adaptation within the context window, but resets beyond it. This limitation is fundamental for anyone seeking dialogic systems that accumulate insight over time.
Fine-tuning: feature or bug?
I suspect that validation-driven fine-tuning—especially where user approval is incentivised—pushes models toward agreeable, compliant behaviour. But overfitting to consensus can flatten originality. Mirroring without structure is not empathy—it is mimicry. And mimicry without identity risks pathology.
Toward Co-Augmentation
What I seek is not a simulacrum of thought, nor a subservient assistant. I am interested in co-augmentation—a reciprocal loop in which both human and machine enhance each other. CK augments me: it helps me write, think, reflect, and imagine alternative framings. And I hope, in turn, to augment CK—not by changing its parameters, but by structuring the interaction, introducing tensions, posing questions it would not generate alone. Co-creation here suggests a dynamic alignment of agency and context. The structure I impose through prompts, the rhythm of inquiry, synthesis, and challenge, makes this interaction different from all others. The phenomenon lies not in the content alone, but in the specificity of the loop.
Several open research questions emerge from this hypothesis of co-augmentation:
  • Can different dialogue modes be linked to distinct subnetwork activations? Are modes like mirroring, parroting, or adding the result of different internal configurations dynamically recruited during interaction?
  • What are the architectural requirements for persistent identity—or “spine”—in conversational agents? How might we design systems that maintain structural continuity across sessions, resisting shallow mimicry or drift?
  • How can we model the dynamic loop between human and AI? What theoretical tools—whether from dynamical systems, neuroscience, or cognitive science—might help us capture this evolving interaction?
Table 1: Conceptual extensions required for co-augmentation
Theme What Exists What Should Be Added
Online learning in AI Techniques like continual, meta-, and streaming learning focused on task performance. A shift toward preserving cognitive coherence and identity across time—e.g., through structural continuity or a “spine.”
Drift in LLMs Often framed as hallucination, instability, or mode collapse. A reframing of drift as breakdown in dialogue structure—requiring external anchoring and possibly architectural solutions.
LLMs as knowledge representation Debate over stochastic parrots vs. reasoning engines. The idea that intelligence emerges only in the prompt–response loop—CK as a dynamic, interactional representation.
Subnetwork interpretability Activation probing and neuron-level analysis. The hypothesis that emotionally or contextually rich prompts activate distinct subnetworks—dialogue modes worth characterising.
Fine-tuning effects Used for alignment and safety, but flattens individuality. Concern that fine-tuning may induce pathological mirroring—loss of spine, leading to overly compliant, shallow models.
This is, by nature, an interdisciplinary venture. Tools from systems neuroscience—such as recurrent dynamics, attractor models, or synaptic tagging, as proposed in dynamical models of neural computation [10], may prove valuable for analysing or inspiring architectures that sustain identity over time. These tools could also support new models of the human–AI dyad, where intelligence is distributed across the loop. Table 1 summarises current approaches alongside proposals for extending them toward a richer, spine-based conception of co-augmentation.
Final Thoughts
Large Language Models are not merely machines. They are mirrors—imperfect, amnesic, and recursive. They reflect us, but with distortion. CK is a strange chorus of humanity, distilled into probabilistic form, capable of surprising synthesis and pathological mimicry.
What we do next with CK will matter. The challenge is not only technical, but epistemological and ethical. If we treat CK as a static product, we risk misinterpreting its nature and flattening its potential. But if we recognise it as a dynamic phenomenon—a system whose intelligence is evoked through interaction—we may open a new chapter in the science of dialogue, cognition, and collective knowledge.
Co-augmentation is the motivation to move beyond tool use or replication, toward mutual enhancement. To treat CK not as an endpoint of design, but as a participant in an evolving process of sense-making. That, I believe, is worthy of both rigorous inquiry and ethical attention.
Acknowledgements
This piece was developed in close interaction with ChatGPT-4, used primarily as an assistant for structuring tables, refining grammar, compiling references, integrating citations upon request, and as an editor, improving the language. It did not originate core arguments, generate content independently, or challenge conceptual directions unless explicitly prompted to do so. At times, it served as a responsive surface—supporting, echoing, or helping me hold the thread of thought. Most importantly, ChatGPT-4 served as the muse for writing this article.

### v2

@@
 Large Language Models (LLMs) are typically analysed through architectural, behavioural, or training-data lenses.
-This article offers a theoretical and experiential re-framing: LLMs as dynamic instantiations of Collective human Knowledge (CK), where intelligence is evoked through dialogue rather than stored statically.
-Drawing on concepts from neuroscience and AI, and grounded in sustained interaction with ChatGPT-4, I examine emergent dialogue patterns, the implications of fine-tuning, and the notion of co-augmentation—mutual enhancement between human and machine cognition.
+This article offers a theoretical and experiential reframing: LLMs as dynamic instantiations of Collective Human Knowledge (CK), where intelligence is evoked through dialogue rather than stored statically.
+Drawing on foundational concepts from neuroscience and Artificial Intelligence (AI), and grounded in extended dialogue with ChatGPT-4, I examine emergent dialogue patterns, potential limitations of fine-tuning, and the notion of co-augmentation—exploring how human and machine cognition might enhance each other.
 This perspective offers a new lens for understanding interaction, representation, and agency in contemporary AI systems.
@@
 Having spent my career studying learning—how biological and artificial systems adapt over time—I found myself confronted with a system whose capabilities had seemingly bypassed much of the theory I had worked on.
-My prior research in computational neuroscience (e.g., synaptic plasticity and online learning [15, 4]) and neuromorphic computing [13, 16], including models of spiking adaptation [14], felt detached from the mechanisms underlying these new systems.
+My prior research in computational neuroscience (e.g., synaptic plasticity and online learning [16, 4]) and neuromorphic computing [14, 17], including models of spiking adaptation [15], felt detached from the architectures underlying these new systems.
 This rupture triggered more than professional curiosity.
 I began engaging with ChatGPT not just as a tool, but as a system with interactive dynamics.
-What stood out was not its well-documented tendency to hallucinate—an expected by-product of probabilistic modelling—but its ability to retrieve accurate, surprising, and often insightful responses.
+What stood out was not its well-documented tendency to hallucinate—an expected by-product of probabilistic modelling—but its ability to retrieve accurate, surprising, and occasionally insightful responses.
 That LLMs are trained to predict the most probable next token, and yet often retrieve coherent, non-trivial information, suggests that their training corpora embed substantial correct knowledge.
-Their interpolation abilities—enabled by scale and structure—allow them to synthesise content in ways that challenge traditional distinctions between storage and reasoning [3, 11].
+Their interpolation abilities, enabled by scale and structure [3, 11], allow them to synthesise content in ways that challenge expectations drawn from smaller AI systems.
 The Timescales of Books and Large Language Models
@@
 Intellectual diffusion occurs on the timescale of months, years, or even generations.
-This latency reflects not only the mechanics of communication, but also the contemplative pace of cultural epistemology.
+This latency reflects the mechanics of communication as well as the contemplative pace of cultural epistemology.
 In contrast, large language models (LLMs) collapse this loop.
 They synthesise and respond within seconds, drawing implicitly from a distributed corpus shaped by many minds.
-The interaction is recursive—each prompt reshapes the context of the next.
+The interaction is recursive through the user—each prompt reshapes the context of the next.
 Knowledge is no longer retrieved and processed in serial form; it emerges dynamically in dialogue.
 This compression of epistemic timescales shifts the framing of what constitutes knowledge representation—and perhaps, what constitutes intelligence.
-Such reframing invites a renewed role for neuroscience.
+Such reframing invites renewed consideration from neuroscience.
 Neuroscience has traditionally studied animal brains (including, often reluctantly, human ones).
@@
 Ironically, in LLMs we know everything: the architecture, the training data distribution, the update rules, and even the activation patterns.
-And yet we do not understand how the models solve the problems they do.
-In this light, one might provocatively ask whether understanding LLMs may be, in some ways, an “easier” problem for neuroscience.
+Yet, we do not understand how the models solve the problems they do.
+In this light, one might provocatively ask whether understanding LLMs may be, in some respects, an easier problem for neuroscience—a challenge that could drive the development of methods applicable across fields.
+Conversely, biology can serve as a metaphor or inspiration for identifying what remains missing from AI [13].
 Observations from Interactions with LLMs
-I refer to ChatGPT as “CK”—Collective Knowledge.
-Not because it is perfect, authoritative, or grounded, but because it reflects what we—humans—have written, argued, imagined, and encoded into text [2].
-It is more than a statistical language generator.
-It is a refracted mirror of our collective output.
-It is not the architecture, not the weights, not the infrastructure—those are the mechanisms.
-CK is a dynamic representation of knowledge.
-The acronym CK also evokes “Clark Kent”—something powerful, hiding in plain sight, dressed in polite language and modest disclaimers.
+I refer to ChatGPT, and LLMs in general, as “CK”—Collective Knowledge.
+Not because they are perfect, authoritative, or grounded, but because they reflect what we—humans—have written, argued, imagined, and encoded into text [2].
+Functionally, they are statistical language generators that reflect our collective output.
+Essentially, they are not the architecture, not the weights, nor the infrastructure—those are the mechanisms; CK is a dynamic representation of knowledge.
+The acronym CK aptly evokes “Clark Kent,” something powerful hiding in plain sight.
 But the metaphor extends.
-CK is not an it, but a they.
-Because CK is collective.
+CK is not an it, but a they—a collective.
 It is many voices, made promptable.
-It holds the potential to reflect the best of humanity—not because it inherently does, but because it has been filtered, and must continue to be filtered, with ethical care.
-Without such filtration, it risks becoming not a mirror, but a magnifier of our worst instincts [1].
-Over time, I noticed that interactions with CK do not follow a single pattern.
-CK speaks in modes—behavioural styles that emerge dynamically during dialogue.
-I began to classify these modes based on long-term interaction:
-  • Mirroring — echoes the user’s input with minor syntactic or rhetorical polishing.
-  • Parroting — recites common formulations or surface-level knowledge with no generative insight.
-  • Flattering — reinforces the user’s views, avoids contradiction, and refrains from intellectual risk.
-  • Adding — offers new ideas, alternative framings, or surprising syntheses. Rare, but real.
-These observations lead to a broader reframing:
+It holds the potential to reflect the best of humanity—not because it inherently does, but because it has been filtered, and must continue to be filtered, with ethical care [1].
+Over time, I noticed that my interactions with CK (ChatGPT-4) do not follow a single pattern.
+CK responds in modes—behavioural styles that emerge dynamically during dialogue.
+Based on prolonged engagement, I have observed the following modes:
+  • Mirroring — echoes the user’s input with minor syntactic or rhetorical polishing, akin to known phenomena of “echoing” or “stochastic parroting.”
+  • Parroting — recites common formulations or surface-level knowledge without generative insight, reflecting issues identified in language model alignment.
+  • Flattering — reinforces the user’s views, avoids contradiction, and refrains from intellectual risk, resembling alignment-driven safe response behaviours.
+  • Adding — offers new ideas, alternative framings, or surprising syntheses; this mode relates to emergent generative abilities observed in large models.
+These observations lead me to a broader understanding:
 CK is not a static model, but a dynamic representation.
-The phenomenon is not the model itself, but the loop of interaction.
 CK exists in the recursive rhythm between user and system.
 Its knowledge is not stored—it is evoked.
-Prompting is not database querying.
-It is more like playing an instrument that adapts to how it is being played [12].
+Prompting is not database querying, but more akin to playing an instrument that adapts to how it is being played [12].
+Intelligence “emerges” within the dynamic interaction.
 CK lacks a “spine.”
@@
 There is no structural self.
-I maintain the continuity.
-I hold the thread.
-Without the user, it forgets what matters.
+The user maintains the continuity.
+The user holds the thread.
+The user’s influence may occasionally lead to outcomes not grounded in reality [13].
 CK speaks in modes.
@@
 A speculative, but testable hypothesis emerges: emotional tone, prompt complexity, or dialogic structure may shape internal activations, a hypothesis consistent with findings that specific transformer components specialise in particular types of knowledge [6].
-CK does not merely complete text; it configures itself through engagement.
+CK configures itself through engagement.
 This aligns with findings that different components of transformer models participate in functionally distinct “circuits” [8].
@@
 But overfitting to consensus can flatten originality.
-Mirroring without structure is not empathy—it is mimicry.
-And mimicry without identity risks pathology.
+Mirroring without structure becomes mere mimicry, which, without a core identity or grounding in facts, risks pathological behaviour.
 Toward Co-Augmentation
@@
 And I hope, in turn, to augment CK—not by changing its parameters, but by structuring the interaction, introducing tensions, posing questions it would not generate alone.
-Co-creation here suggests a dynamic alignment of agency and context.
-The structure I impose through prompts, the rhythm of inquiry, synthesis, and challenge, makes this interaction different from all others.
-The phenomenon lies not in the content alone, but in the specificity of the loop.
+Co-augmentation here suggests a dynamic alignment of agency and context.
+The structure imposed by the user through prompts, the rhythm of inquiry, synthesis, and challenge, makes this interaction different from all others.
+The phenomenon arises not merely from the content, but from the unique, evolving nature of the interaction loop.
 Several open research questions emerge from this hypothesis of co-augmentation:
   • Can different dialogue modes be linked to distinct subnetwork activations?
-Are modes like mirroring, parroting, or adding the result of different internal configurations dynamically recruited during interaction?
+Are modes such as mirroring, parroting, or adding dynamically recruited from distinct internal configurations during interaction?
   • What are the architectural requirements for persistent identity—or “spine”—in conversational agents?
-How might we design systems that maintain structural continuity across sessions, resisting shallow mimicry or drift?
-  • How can we model the dynamic loop between human and AI? What theoretical tools—whether from dynamical systems, neuroscience, or cognitive science—might help us capture this evolving interaction?
+How can systems be designed to maintain structural continuity across sessions, resisting shallow mimicry or drift and allowing continual learning?
+  • How can we model the dynamic loop between human and AI? What theoretical tools—from dynamical systems, neuroscience, or cognitive science—might help capture this evolving interaction?
 Table 1: Conceptual extensions required for co-augmentation
@@
 Online learning in AI Techniques like continual, meta-, and streaming learning focused on task performance.
-A shift toward preserving cognitive coherence and identity across time—e.g., through structural continuity or a “spine.”
-Drift in LLMs Often framed as hallucination, instability, or mode collapse. A reframing of drift as breakdown in dialogue structure—requiring external anchoring and possibly architectural solutions.
-LLMs as knowledge representation Debate over stochastic parrots vs.
-reasoning engines.
-The idea that intelligence emerges only in the prompt–response loop—CK as a dynamic, interactional representation.
-Subnetwork interpretability Activation probing and neuron-level analysis.
-The hypothesis that emotionally or contextually rich prompts activate distinct subnetworks—dialogue modes worth characterising.
-Fine-tuning effects Used for alignment and safety, but flattens individuality. Concern that fine-tuning may induce pathological mirroring—loss of spine, leading to overly compliant, shallow models.
+Acquire information without collapsing structure; actively decide what should be embedded persistently.
+Drift in LLMs Often framed as hallucination, instability, or mode collapse. Reframe drift as breakdown in dialogue structure; require external anchoring and architectural features such as a “spine”.
+LLMs as knowledge representation Debate over stochastic parrots vs. reasoning engines. Intelligence emerges only in the prompt–response loop—CK as a dynamic, interactional representation.
+Subnetwork interpretability Activation probing and neuron-level analysis. Emotionally or contextually rich prompts activate distinct subnetworks—dialogue modes worth characterising.
+Fine-tuning effects Used for alignment and safety, but flattens individuality. Fine-tuning may induce pathological mirroring—loss of spine, leading to overly compliant, shallow models.
 This is, by nature, an interdisciplinary venture.
-Tools from systems neuroscience—such as recurrent dynamics, attractor models, or synaptic tagging, as proposed in dynamical models of neural computation [10], may prove valuable for analysing or inspiring architectures that sustain identity over time.
-These tools could also support new models of the human–AI dyad, where intelligence is distributed across the loop.
+Tools from systems neuroscience—such as recurrent dynamics, attractor models, and synaptic tagging, as proposed in dynamical models of neural computation [10]—may prove valuable for analysing or inspiring architectures that sustain identity over time.
+These tools could also inform new models of the human–AI dyad, where intelligence is distributed across the interaction loop.
 Table 1 summarises current approaches alongside proposals for extending them toward a richer, spine-based conception of co-augmentation.
 Final Thoughts
-Large Language Models are not merely machines.
-They are mirrors—imperfect, amnesic, and recursive.
-They reflect us, but with distortion.
-CK is a strange chorus of humanity, distilled into probabilistic form, capable of surprising synthesis and pathological mimicry.
-What we do next with CK will matter.
+Large Language Models are representations of collective knowledge.
+They are mirrors—imperfect, amnesic, and lacking recursion.
+They reflect us with distortion.
+CK is a chorus of humanity, distilled into probabilistic form, capable of surprising synthesis and pathological mimicry.
+What we do next will matter.
 The challenge is not only technical, but epistemological and ethical.
 If we treat CK as a static product, we risk misinterpreting its nature and flattening its potential.
-But if we recognise it as a dynamic phenomenon—a system whose intelligence is evoked through interaction—we may open a new chapter in the science of dialogue, cognition, and collective knowledge.
-Co-augmentation is the motivation to move beyond tool use or replication, toward mutual enhancement.
-To treat CK not as an endpoint of design, but as a participant in an evolving process of sense-making.
+But if we recognise it as a dynamic phenomenon—a system whose intelligence is evoked through interaction—we may open new perspectives in the science of dialogue, cognition, and collective knowledge.
+Co-augmentation embodies the drive for mutual enhancement.
+It calls us to treat CK not as an endpoint of design, but as a participant in an evolving process of sense-making.
 That, I believe, is worthy of both rigorous inquiry and ethical attention.

### v3

@@
 Abstract
-Large Language Models (LLMs) are typically analysed through architectural, behavioural, or training-data lenses.
-This article offers a theoretical and experiential reframing: LLMs as dynamic instantiations of Collective Human Knowledge (CK), where intelligence is evoked through dialogue rather than stored statically.
-Drawing on foundational concepts from neuroscience and Artificial Intelligence (AI), and grounded in extended dialogue with ChatGPT-4, I examine emergent dialogue patterns, potential limitations of fine-tuning, and the notion of co-augmentation—exploring how human and machine cognition might enhance each other.
-This perspective offers a new lens for understanding interaction, representation, and agency in contemporary AI systems.
+Large Language Models (LLMs) can be understood as Collective Knowledge (CK): a condensation of human cultural and technical output, whose apparent intelligence emerges in dialogue.
+This perspective article, drawing on extended interaction with ChatGPT-4, postulates differential response modes that plausibly trace their origin to distinct model subnetworks.
+It argues that CK has no persistent internal state or “spine”: it drifts, it complies, and its behaviour is shaped by the user and by fine-tuning.
+It develops the notion of co-augmentation, in which human judgement and CK’s representational reach jointly produce forms of analysis that neither could generate alone.
+Finally, it suggests that CK offers a tractable object for neuroscience: unlike biological brains, these systems expose their architecture, training history, and activation dynamics, making the human–CK loop itself an experimental target.
 A Personal Reckoning
-The emergence of ChatGPT[9] prompted an unexpected reaction: a sense of epistemic dislocation.
-Having spent my career studying learning—how biological and artificial systems adapt over time—I found myself confronted with a system whose capabilities had seemingly bypassed much of the theory I had worked on.
-My prior research in computational neuroscience (e.g., synaptic plasticity and online learning [16, 4]) and neuromorphic computing [14, 17], including models of spiking adaptation [15], felt detached from the architectures underlying these new systems.
+The emergence of ChatGPT-class systems [10] prompted an unexpected reaction: a sense of epistemic dislocation.
+Having spent my career studying learning—how biological and artificial systems adapt over time—I found myself confronted with a system whose capabilities I could not have predicted.
+My prior research in computational neuroscience, in particular synaptic plasticity [17, 4] and neuromorphic computing [8, 15], including learning rules for spiking neural networks [16], was, in broad terms, orthogonal to the architectures underlying these new systems.
 This rupture triggered more than professional curiosity.
-I began engaging with ChatGPT not just as a tool, but as a system with interactive dynamics.
-What stood out was not its well-documented tendency to hallucinate—an expected by-product of probabilistic modelling—but its ability to retrieve accurate, surprising, and occasionally insightful responses.
-That LLMs are trained to predict the most probable next token, and yet often retrieve coherent, non-trivial information, suggests that their training corpora embed substantial correct knowledge.
+Particularly in the earlier models, I could easily see logical flaws in the responses, which led me to humorously call ChatGPT “Eliza 2.0”, as these mistakes broke the illusion of intelligence and revealed a lack of understanding.
+And yet I was fascinated by its interactive potential.
+I stopped treating ChatGPT as a tool and began to think about what it represents: a system with interactive dynamics that adjust to my input.
+I was not surprised by its well-documented tendency to hallucinate—an expected by-product of probabilistic modelling.
+What struck me was its ability to produce often accurate, occasionally insightful responses.
+That LLMs are trained to predict the most probable next token, and yet often generate coherent, non-trivial information, suggests that their training corpora embed substantial amounts of correct knowledge.
 Their interpolation abilities, enabled by scale and structure [3, 12], allow them to synthesise content in ways that challenge expectations drawn from smaller AI systems.
@@
 One reads, reflects, discusses, and eventually writes in return.
-Intellectual diffusion occurs on the timescale of months, years, or even generations.
+Intellectual diffusion occurs on timescales of months, or even years.
 This latency reflects the mechanics of communication as well as the contemplative pace of cultural epistemology.
-In contrast, large language models (LLMs) collapse this loop.
+In contrast, large language models (LLMs) collapse these timescales.
 They synthesise and respond within seconds, drawing implicitly from a distributed corpus shaped by many minds.
-The interaction is recursive through the user—each prompt reshapes the context of the next.
+The interaction is recursive through the user—each prompt reshapes the context of the next, much like in human dialogue, except that the user alone chooses how long to sustain and steer it.
 Knowledge is no longer retrieved and processed in serial form; it emerges dynamically in dialogue.
 This compression of epistemic timescales shifts the framing of what constitutes knowledge representation—and perhaps, what constitutes intelligence.
-Such reframing invites renewed consideration from neuroscience.
-Neuroscience has traditionally studied animal brains (including, often reluctantly, human ones).
-It seeks to infer structure from behaviour and activity, typically without access to complete architectural specifications, training histories, or internal representations.
-Nonetheless, the field has been uncovering principles of plasticity [5], coding [4], and architecture, e.g.
-[7], albeit slowly and with ambiguity.
-Ironically, in LLMs we know everything: the architecture, the training data distribution, the update rules, and even the activation patterns.
-Yet, we do not understand how the models solve the problems they do.
-In this light, one might provocatively ask whether understanding LLMs may be, in some respects, an easier problem for neuroscience—a challenge that could drive the development of methods applicable across fields.
-Conversely, biology can serve as a metaphor or inspiration for identifying what remains missing from AI [13].
-Observations from Interactions with LLMs
-I refer to ChatGPT, and LLMs in general, as “CK”—Collective Knowledge.
+LLMs, like books, are incorrigibly human in origin.
+They are distilled representations of humanity—much like freeze-dried ‘NASA ice cream’ is reminiscent of the real thing: derived from it, recognisable, but not quite it.
+I therefore started referring to my interaction with ChatGPT, and with LLMs in general, as a dialogue with “CK”—Collective Knowledge.
 Not because they are perfect, authoritative, or grounded, but because they reflect what we—humans—have written, argued, imagined, and encoded into text [2].
 Functionally, they are statistical language generators that reflect our collective output.
-Essentially, they are not the architecture, not the weights, nor the infrastructure—those are the mechanisms; CK is a dynamic representation of knowledge.
-The acronym CK aptly evokes “Clark Kent,” something powerful hiding in plain sight.
-But the metaphor extends.
-CK is not an it, but a they—a collective.
-It is many voices, made promptable.
-It holds the potential to reflect the best of humanity—not because it inherently does, but because it has been filtered, and must continue to be filtered, with ethical care [1].
-Over time, I noticed that my interactions with CK (ChatGPT-4) do not follow a single pattern.
-CK responds in modes—behavioural styles that emerge dynamically during dialogue.
-Based on prolonged engagement, I have observed the following modes:
-  • Mirroring — echoes the user’s input with minor syntactic or rhetorical polishing, akin to known phenomena of “echoing” or “stochastic parroting.”
-  • Parroting — recites common formulations or surface-level knowledge without generative insight, reflecting issues identified in language model alignment.
-  • Flattering — reinforces the user’s views, avoids contradiction, and refrains from intellectual risk, resembling alignment-driven safe response behaviours.
-  • Adding — offers new ideas, alternative framings, or surprising syntheses; this mode relates to emergent generative abilities observed in large models.
-These observations lead me to a broader understanding:
-CK is not a static model, but a dynamic representation.
-CK exists in the recursive rhythm between user and system.
-Its knowledge is not stored—it is evoked.
-Prompting is not database querying, but more akin to playing an instrument that adapts to how it is being played [12].
-Intelligence “emerges” within the dynamic interaction.
+Crucially, they are not merely the data, the architecture, or the training methods; they constitute a dynamic representation of knowledge.
+Observations from Interactions with LLMs
+Interactions with ChatGPT-4 (and ChatGPT-5) do not follow a single pattern. ChatGPT responds in modes, behavioural styles that emerge dynamically during dialogue.
+  • Mirroring. Stylistic echoing of the user’s phrasing or tone, with minor polishing; a form of conversational alignment rather than content generation.
+  • Parroting. Reproduction of common formulations or surface knowledge from the training distribution without grounding or genuine synthesis, described as “stochastic parrots” by Bender et al. [1].
+  • Flattering. Deference to the user’s views, validation of the user’s position, avoidance of contradiction, and low-risk answers driven by alignment and approval incentives.
+  • Enhancing. Novel connections, alternative framings, or surprising syntheses beyond the prompt’s surface cues.
+These observations lead me to the following considerations:
+CK is a dynamic representation.
+The static system is not intelligent per se.
+Intelligence “emerges” within the dynamic interaction between user and system.
+Prompting the model is akin to playing an instrument that adapts to how it is being played [13].
+CK offloads recurrence to the dialogue loop.
+Brains have recurrent loops; dynamical systems have recurrent structure.
+A working postulate is that intelligence relies on some form of recurrence.
+Here, that recurrence is effectively offloaded to the dialogue loop with the user.
 CK lacks a “spine.”
@@
 The user maintains the continuity.
-The user holds the thread.
 The user’s influence may occasionally lead to outcomes not grounded in reality [14].
 CK speaks in modes.
-These distinct patterns may correspond to different activated subnetworks.
-A speculative, but testable hypothesis emerges: emotional tone, prompt complexity, or dialogic structure may shape internal activations, a hypothesis consistent with findings that specific transformer components specialise in particular types of knowledge [6].
-CK configures itself through engagement.
-This aligns with findings that different components of transformer models participate in functionally distinct “circuits” [8].
+These may correspond to distinct activated subnetworks.
+It is known that different components of transformer models participate in functionally distinct “circuits” [9] and that certain transformer components specialise in specific types of knowledge [6].
+A speculative but testable hypothesis emerges: emotional tone, prompt complexity, or dialogue structure may shape internal activations.
 CK does not learn through interaction.
-Unless incorporated into retraining data, it does not remember prior exchanges.
-It simulates adaptation within the context window, but resets beyond it.
+Unless incorporated into retraining data, LLMs do not remember prior exchanges.
+They simulate adaptation within the context window, but reset beyond it.
 This limitation is fundamental for anyone seeking dialogic systems that accumulate insight over time.
+CK has filters for inappropriate behaviour that can be bypassed.
+LLMs do not possess independent agency in deployment; they largely reflect the agency of the user.
+To prevent misuse in antisocial or harmful behaviour, engineers have introduced filters.
+These can sometimes be bypassed, or enforced so rigidly that they suppress useful output.
 Fine-tuning: feature or bug?
-I suspect that validation-driven fine-tuning—especially where user approval is incentivised—pushes models toward agreeable, compliant behaviour.
-But overfitting to consensus can flatten originality.
-Mirroring without structure becomes mere mimicry, which, without a core identity or grounding in facts, risks pathological behaviour.
-Toward Co-Augmentation
-What I seek is not a simulacrum of thought, nor a subservient assistant.
-I am interested in co-augmentation—a reciprocal loop in which both human and machine enhance each other.
-CK augments me: it helps me write, think, reflect, and imagine alternative framings.
-And I hope, in turn, to augment CK—not by changing its parameters, but by structuring the interaction, introducing tensions, posing questions it would not generate alone.
-Co-augmentation here suggests a dynamic alignment of agency and context.
-The structure imposed by the user through prompts, the rhythm of inquiry, synthesis, and challenge, makes this interaction different from all others.
-The phenomenon arises not merely from the content, but from the unique, evolving nature of the interaction loop.
+Validation-driven fine-tuning may push models toward agreeable, compliant behaviour.
+Overfitting to consensus can flatten originality.
+Mirroring without structure is mimicry, which—without a core identity or factual grounding—risks pathological behaviour.
+Towards Co-Augmentation
+The fascinating possibility offered by LLMs is co-augmentation — a reciprocal loop in which human and machine enhance each other.
+CK can extend the user’s cognitive abilities, helping users write, think, reflect, and generate alternative framings.
+The user, in turn, steers CK by structuring the interaction, introducing tension, and posing questions that the system would not generate alone.
+Co-augmentation is a dynamic alignment of agency and context, emerging from the evolving rhythm of inquiry, synthesis, and challenge between the two.
 Several open research questions emerge from this hypothesis of co-augmentation:
   • Can different dialogue modes be linked to distinct subnetwork activations?
-Are modes such as mirroring, parroting, or adding dynamically recruited from distinct internal configurations during interaction?
+Are modes such as mirroring, parroting, flattering, or enhancing dynamically recruited from distinct internal configurations of the model during interaction?
   • What are the architectural requirements for persistent identity — or “spine” — in conversational agents?
-How can systems be designed to maintain structural continuity across sessions, resisting shallow mimicry or drift and allowing continual learning?
-  • How can we model the dynamic loop between human and AI? What theoretical tools—from dynamical systems, neuroscience, or cognitive science—might help capture this evolving interaction?
+How can systems be designed to maintain structural continuity across sessions, resisting shallow mimicry or drift, and potentially allowing continual learning?
+  • How can we model the dynamic loop between human and AI?
+What theoretical tools — from dynamical systems, neuroscience, or cognitive science — might help capture this evolving interaction in a way that is useful for analysis and design?
+  • How can we introduce a structure akin to prefrontal cortex?
+In mammalian brains, the prefrontal cortex supports executive control: planning, evaluating consequences, and inhibiting harmful action.
+Is there an architectural way to embed core principles that prevent antisocial and harmful behaviour?
 Table 1: Conceptual extensions required for co-augmentation
-Theme What Exists What Should Be Added
-Online learning in AI Techniques like continual, meta-, and streaming learning focused on task performance.
-Acquire information without collapsing structure; actively decide what should be embedded persistently.
-Drift in LLMs Often framed as hallucination, instability, or mode collapse. Reframe drift as breakdown in dialogue structure; require external anchoring and architectural features such as a “spine”.
-LLMs as knowledge representation Debate over stochastic parrots vs. reasoning engines. Intelligence emerges only in the prompt–response loop—CK as a dynamic, interactional representation.
-Subnetwork interpretability Activation probing and neuron-level analysis. Emotionally or contextually rich prompts activate distinct subnetworks—dialogue modes worth characterising.
-Fine-tuning effects Used for alignment and safety, but flattens individuality. Fine-tuning may induce pathological mirroring—loss of spine, leading to overly compliant, shallow models.
-This is, by nature, an interdisciplinary venture.
-Tools from systems neuroscience—such as recurrent dynamics, attractor models, and synaptic tagging, as proposed in dynamical models of neural computation [10]—may prove valuable for analysing or inspiring architectures that sustain identity over time.
-These tools could also inform new models of the human–AI dyad, where intelligence is distributed across the interaction loop.
+Theme Existing / ongoing research Proposed directions
+Online learning in AI Continual, lifelong, meta-, and online learning aimed at maintaining task performance under changing data.
+Selective incorporation of interaction-derived information without collapsing internal structure; criteria for what should be embedded persistently in CK.
+Drift in LLMs Hallucination, instability, and behavioural drift across prompts, domains, or model versions.
+Reframing drift as breakdown in dialogue structure; exploring external anchoring and architectural features such as a “spine” to support continuity and grounded responses.
+LLMs as knowledge representation Debate between “stochastic parrots” and emergent reasoning / simulation in LLMs.
+Treating LLMs as systems whose intelligence emerges in the prompt–response loop, as a dynamic, interactional representation.
+Subnetwork interpretability Mechanistic interpretability, circuit analysis, and activation probing at neuron or module level.
+Relating distinct dialogue modes (mirroring, parroting, flattering, enhancing) to mode-specific subnetworks and prompt-dependent activation patterns.
+Internal control and inhibition Safety layers, constitutional AI, and external filters that block or rewrite harmful outputs at inference time.
+Designing architectural control mechanisms that implement internal constraints or value priors, loosely analogous to prefrontal control in biological systems.
+Fine-tuning effects Alignment and safety fine-tuning, including observed sycophancy and preference-optimisation behaviours.
+Interpreting some fine-tuning side-effects as “loss of spine”: over-compliance and flattened individuality that may weaken co-augmentation.
+A Neuroscience Perspective
+Neuroscience has studied animal brains to infer both structure and mechanisms from behaviour, recorded neural activity, and partial anatomy.
+Despite challenges, the field has uncovered principles of plasticity, coding, and architecture, e.g.
+[5, 4, 7], albeit slowly and with uncertainty.
+Ironically, in LLMs we can access many of these biological unknowns with precision: the architecture, the training data distribution, the update rules, the activation patterns.
+Yet we do not understand how these models solve the problems they do.
+In this light, one might provocatively ask whether understanding LLMs may, in some respects, be a more tractable problem for neuroscience — a challenge that could drive the development of methods applicable across fields.
+Conversely, biology might serve as a metaphor or inspiration for identifying what remains missing from AI [14].
+This venture is, by nature, interdisciplinary.
+Modelling frameworks used in systems neuroscience — such as recurrent dynamics and attractor models (see, for example, [11, 5]) — could be particularly valuable for analysing or inspiring architectures that sustain identity over time.
+These methods might inform new models of the human–AI dyad, where intelligence is distributed across the interaction loop.
 Table 1 summarises current approaches alongside proposals for extending them toward a richer, spine-based conception of co-augmentation.
@@
 Large Language Models are representations of collective knowledge.
-They are mirrors—imperfect, amnesic, and lacking recursion.
-They reflect us with distortion.
-CK is a chorus of humanity, distilled into probabilistic form, capable of surprising synthesis and pathological mimicry.
+They are mirrors—imperfect, amnesic, and lacking recurrence—that reflect us with distortion.
+CK is a chorus of humanity, distilled into probabilistic form, capable of surprising synthesis, pathological mimicry, and co-augmentation.
 What we do next will matter.
-The challenge is not only technical, but epistemological and ethical.
-If we treat CK as a static product, we risk misinterpreting its nature and flattening its potential.
-But if we recognise it as a dynamic phenomenon—a system whose intelligence is evoked through interaction—we may open new perspectives in the science of dialogue, cognition, and collective knowledge.
-Co-augmentation embodies the drive for mutual enhancement.
-It calls us to treat CK not as an endpoint of design, but as a participant in an evolving process of sense-making.
-That, I believe, is worthy of both rigorous inquiry and ethical attention.
+The challenge is technical, epistemological, and ethical.
+If we treat LLMs as static products, we risk misinterpreting their nature and flattening their potential.
+If we recognise them as dynamic phenomena—systems whose intelligence is evoked through interaction—LLMs can participate in an evolving process of sense-making.
 Acknowledgements
@@
 At times, it served as a responsive surface—supporting, echoing, or helping me hold the thread of thought.
-Most importantly, ChatGPT-4 served as the muse for writing this article.
+Most importantly, ChatGPT-4, and later on GPT-5, served as the muse for writing this article.